OpenStudy (anonymous):

uhh, i dont know any other way to put this up here, to find to the value of x and y: 5x-4y=3y-x+8

OpenStudy (bahrom7893):

that's 1 equation and 2 unknowns.. not enough info

OpenStudy (anonymous):

hey let me post a pic of it for u you will better understand i hope! :D

OpenStudy (aripotta):

that might help

OpenStudy (anonymous):

here yer go!! this problem is evil .

OpenStudy (aripotta):

hm.

OpenStudy (bahrom7893):

x+y=2x-4 5x-4y=3y-x+8

OpenStudy (bahrom7893):

Solve the 2 equations

OpenStudy (anonymous):

Important details are important, yes they are. Especially if they are geometry details.

OpenStudy (bahrom7893):

ok im going to bed folks.. can someone help him/her solve the equations?

OpenStudy (anonymous):

Sure, I got this. I think.

OpenStudy (aripotta):

oooh i got it. omg. i feel smart

OpenStudy (anonymous):

x-2x+4=y -x+4=y -7y=-6x+8 7y=6x-8 Plug in: 7(4-x)=6x-8 28-7x=6x-8 36=13x x=13/36

OpenStudy (anonymous):

Part two...

OpenStudy (anonymous):

-13/36+4=y y=131/36

OpenStudy (aripotta):

that doesn't look right to me :/

OpenStudy (anonymous):

No? It's late here, where did I mess up? :\

OpenStudy (anonymous):

just think because it is a regular octagon all sides are equal so all equations are equal, Choose an equation that contains one variable, solve for it and substitute it into another equation.

OpenStudy (anonymous):

But they all have both x and y in them, so you have to make a system of equations like bahrom7893 did.

OpenStudy (tyteen4a03):

Remember to change the subject to help you ;)

OpenStudy (aripotta):

bahrom said to solve these: x+y=2x-4, y = x - 4 5x-4y=3y-x+8, 5x - 4(x - 4) = 3(x - 4) - x + 8 5x - 4x + 16 = 3x - 12 - x + 8 x + 16 = 2x - 4 x + 20 = 2x 20 = x

OpenStudy (anonymous):

So I made stupid algebra mistakes then. Right, makes sense, thank you very kindly! XD

OpenStudy (anonymous):

you can still solve for one variable and substitute it into another equation to make that equation only have one variable. Then you can get a numerical value for that variable. You will then have to take this value and plug it into an equation to get the other.

OpenStudy (aripotta):

y would = 16

OpenStudy (aripotta):

then for the side that's x + y...20 + 16 = 36 and because there's 8 sides...36 x 8 = 288

OpenStudy (aripotta):

so 288 would be the perimeter....

OpenStudy (anonymous):

thanks so much guys! you all make so sexy to be learning!!

OpenStudy (aripotta):

lol

OpenStudy (anonymous):

Learning is always sexy.

OpenStudy (anonymous):

sexy to be living in USA

OpenStudy (anonymous):

anyway once again thanks, its 2am, shuteye is needed

OpenStudy (anonymous):

Good night and good luck.

hartnn (hartnn):

good nite, have sexy dreams....